OpenStudy (anonymous):

what is an equation of the line that has slope -4 and passes through the point (-2,-5) A. y= -4x-8 B. y= -4x-13 c. y=-4x-5 D. y=-4x+3

OpenStudy (anonymous):

y = mx + b m = -4 (-2,-5) x = -2 and y = -5 now we sub y = mx + b -5 = -4(-2) + b -5 = 8 + b -5 - 8 = b - 13 = b your equation is : y = -4x - 13

OpenStudy (anonymous):

thanks